# Z-Dip: a validated generalization of the Dip Test

### v1

@@ Abstract
Detecting multimodality in empirical distributions is a fundamental problem in statistics and data analysis, with applications ranging from clustering to social science. Hartigan’s Dip Test is a classical nonparametric procedure for testing unimodality versus multimodality, but its interpretation is hindered by strong dependence on sample size and the need for lookup tables. We introduce the Z-Dip, a standardized extension of the Dip Test that removes sample-size dependence by comparing observed Dip values to simulated null distributions. We calibrate a universal decision threshold for Z-Dip via simulation and bootstrap resampling, providing a unified criterion for multimodality detection. In the final section, we also propose a downsampling-based approach to further mitigate residual sample-size effects in very large datasets. Lookup tables and software implementations are made available for efficient use in practice.
@@ 1 Introduction
The ability to test whether a distribution is unimodal or multimodal is crucial across many domains. In clustering, multimodality indicates the presence of distinct groups. In social and political science, multimodality in opinion distributions is often interpreted as evidence of polarization. Statistical tests of modality thus provide an essential link between empirical data and theoretical interpretation.
Hartigan and Hartigan’s Dip Test [1] remains one of the most widely used nonparametric methods for testing unimodality, and has found applications across diverse fields [2, 3, 4]. The test statistic, known as the Dip statistic, measures the maximum deviation between the empirical distribution function and the closest unimodal distribution function. Critical values and $p$-values under the null hypothesis of unimodality are determined by simulation under a uniform distribution.
Despite its popularity, the Dip Test suffers from three major limitations. First, its critical values are strongly dependent on sample size, requiring large lookup tables or resimulation for each application. Second, the test statistic is not standardized, preventing meaningful comparisons of the strength of multimodality across datasets. Third, for large samples, the Dip Test tends to reject unimodality even for nearly unimodal distributions, leading to an undesirable loss of resolution.
In this paper, we propose a simple yet effective generalization of the Dip Test, which we call the Z-Dip. By standardizing the observed Dip against its null distribution, the Z-Dip provides a scale-free measure of deviation from unimodality that can be compared across sample sizes. Moreover, we calibrate a single universal threshold for statistical significance, eliminating the need for sample-size–specific lookup tables. The resulting algorithm, implemented in R and distributed alongside this paper, retains the same $O(n\log n)$ computational complexity as the original Dip Test while offering more consistent and interpretable results.
We validate the Z-Dip both on simulated samples, and on an empirical sample of more than 117,000 opinion distributions, finding effectively complete agreement between the proposed Z-Dip threshold and the Dip Test’s $p$-value decisions. Finally, we address the issue of information loss in large samples and introduce an alternative correction, which enables consistent comparisons across large datasets.
By explicitly characterizing how sample size affects standardized multimodality statistics, and by offering a transparent correction procedure, this work aims to improve the interpretability and comparability of such metrics. The correction can be implemented with minimal changes to existing workflows and with virtually no added computational overhead. In this sense, the simplicity of the approach is its main strength: it provides a clear and general remedy to a problem that is otherwise difficult to detect.
@@ 2 Background
2.1 Hartigan’s Dip Test
Hartigan’s Dip Test, introduced by Hartigan and Hartigan in 1985 [1], is a classical nonparametric method for testing whether a sample distribution is unimodal or multimodal. Formally, it tests the null hypothesis of unimodality against the alternative of multimodality, using the uniform distribution as the “least favorable unimodal null”. The test quantifies how far the empirical cumulative distribution function (ECDF) of the data deviates from its respective closest unimodal cumulative distribution function: intuitively, a unimodal distribution is one whose density has a single peak, which implies that its cumulative distribution is convex up to the mode and concave thereafter. The Dip Test exploits this shape property by finding the unimodal CDF that minimizes the maximum vertical distance from the observed ECDF. This minimum distance, known as the Dip statistic, summarizes how strongly the observed distribution departs from unimodality. In essence, the Dip measures the smallest possible maximum discrepancy between the observed data and any unimodal distribution. Larger Dip values indicate greater deviation from unimodality, and therefore stronger evidence of multimodality.
The sampling distribution of the Dip under the null hypothesis is obtained by simulation from the uniform distribution on $[0,1]$, which represents the boundary case of a unimodal distribution. Hartigan and Hartigan provided tabulated critical values for a range of sample sizes, allowing the test to be implemented via lookup or resimulation. A key strength of the Dip Test lies in its simplicity: it is fully nonparametric and requires no user-specified parameters, bandwidths, or smoothing assumptions. Unlike kernel- or mixture-based methods, the Dip Test can be applied “out of the box,” with no tuning or model selection. This property has contributed to its wide adoption, as researchers can directly apply it to empirical data without additional calibration or subjective choices. Thanks to its aforementioned simplicity and robustness, since its introduction, the Dip Test has been widely adopted across fields [5, 2, 3, 4].
@@ 2.2 Limitations of the Classical Dip Test
    Dependence on sample size. The null distribution of the Dip statistic varies strongly with sample size $N$. As $N$ increases, the Dip values expected under unimodality decrease systematically, since larger samples provide finer-grained approximations to the true distribution. Consequently, each sample size requires its own set of critical values or a separate Monte Carlo calibration. This dependence complicates both interpretation and automation, as results from samples of different sizes cannot be directly compared.
@@
    Lack of standardization. The Dip statistic itself is not standardized. Its raw magnitude depends on the scale of sampling variability and on $N$, making it unsuitable for quantitative comparison across datasets. For example, a Dip of $0.1$ may signal multimodality in a small sample, but be negligible in a large one. Without a standardized scale, the test can only yield binary decisions (reject or not reject unimodality), rather than interpretable effect sizes that quantify “how multimodal” a distribution is. However, the Dip statistic already encapsulates this information: it measures the strength of deviation from unimodality, but this signal is obscured by sample-size dependence.
@@
    Inconsistencies at large sample sizes. At large $N$, the Dip Test becomes increasingly sensitive to even minor deviations from the ideal unimodal shape. Because the null distribution of the Dip shrinks toward zero as $N$ grows, the smallest irregularities, such as sampling noise or slight skewness, are sufficient to reject unimodality. In practice, this produces false positives in large samples: distributions that are visually unimodal but statistically classified as multimodal. This issue means that the Dip Test’s binary decision rule becomes less informative with increasing $N$, effectively turning every large sample into a “significant” case. As a result, the test loses its ability to distinguish meaningful multimodality (e.g., distinct peaks) from trivial numerical deviations.
Together, these limitations point to the need for a standardized, scale-free reformulation of the Dip Test that preserves its parameter-free simplicity while allowing meaningful comparison across sample sizes, maintaining interpretability at large $N$, and eliminating the need for case-specific calibration. The next section introduces the Z-Dip, which addresses these issues by standardizing the Dip relative to its null distribution.
@@ 3.1 Definition of the Z-Dip Statistic
We propose a standardized version of Hartigan’s Dip statistic, which we call the Z-Dip. For an observed dataset of size $N$, let $Dip_{obs}$ denote its Dip statistic’s value. We define
$Z\text{-}Dip=z=\frac{Dip_{obs}-\mu_{N}}{\sigma_{N}},$ (1)
@@
where $\mu_{N}$ and $\sigma_{N}$ denote the mean and standard deviation of Dip values under the null distribution of $N$-sized samples drawn from the uniform distribution on $[0,1]$.
The score proposed in Formula 1 quantifies how strongly the observed distribution deviates from unimodality, in units of standard deviation relative to the null expectation. Positive values indicate stronger evidence of multimodality, while values near zero or negative indicate conformity with unimodality. Because it rescales the Dip relative to its null, the Z-Dip provides a scale-free, comparable measure across different sample sizes.

### v2

# Z-Dip: a standardized measure for data modality assessment

@@ Abstract
Detecting multimodality in empirical distributions is a fundamental problem in statistics and data analysis, with applications ranging from clustering to the study of complex systems. In practice, however, assessing departures from unimodality in a consistent and comparable way remains challenging. Widely used methods such as Hartigan and Hartigan’s Dip Test illustrate these difficulties, as the interpretation of their statistics depends strongly on sample size, requires calibration to determine significance, and, for large samples, exhibit increasing sensitivity, leading to rejection of unimodality for arbitrarily small deviations from the null. We introduce Z-Dip, a standardized measure of multimodality that addresses these limitations. By treating the Dip statistic as a random variable under the null hypothesis of unimodality and standardizing its observed value, the proposed approach yields scores that are directly comparable across datasets of different sizes. Using simulation-based calibration, we derive a universal decision threshold that closely reproduces classical Dip Test decisions without requiring sample-size-specific adjustments. Extensive validation on simulated data and on more than 88,000 empirical opinion distributions shows near-perfect agreement with the classical Dip Test while providing a more interpretable and comparable measure of modality. Finally, we propose a downsampling-based correction that mitigates residual sensitivity in extremely large samples. Open-source software and reference tables are provided to facilitate practical adoption.
keywords:
Modality assessment, Distributional shape analysis, Nonparametric statistics
journal: .
[…]
[inst1]organization=Department of Social Sciences and Economics, Sapienza University of Rome, addressline=P.le Aldo Moro, 5, city=Rome, postcode=00185, state=Lazio, country=Italy
[…]
[inst2]organization=Department of Computer Science, Sapienza University of Rome, addressline=Viale Regina Elena, 295, city=Rome, postcode=00161, state=Lazio, country=Italy
@@ 1 Introduction
Multimodality is a relevant feature of many empirical distributions, reflecting the presence of multiple concentrations of data around specific values, which are often indicative of heterogeneous generating processes or coexisting regimes underlying the observed phenomenon. Correctly identifying departures from unimodality is therefore of broad interest across empirical sciences, both as a descriptive task and as a way to detect nontrivial structure in observed data: as modern applications increasingly rely on large-scale, high-dimensional, and weakly structured data, automated and reliable methods for detecting and quantifying departures from unimodality have become essential to inform decisions [1, 2]. Examples of multimodal patterns arise across a wide range of domains. In complex networks, the modality of degree distributions informs structural inference [3], distinguishing between homogeneous connectivity and more structured organizations such as core–periphery architectures [4, 5, 6]. Multimodality has also been observed in other graph-related quantities, including path-length statistics [7] and redundancy measures [8]. Similar considerations apply to social systems, where multimodal distributions of beliefs or stances emerge from interaction and aggregation processes. In this context, bimodality is often used as a proxy for political polarization [9], and has been linked to fragmented discourse and limited cross-group interaction [10, 11, 12].
Following this perspective, multimodality is not merely a descriptive property of empirical data, but can also be an indicator of structural properties of the system or a description of complex underlying processes. Correctly identifying it, particularly across large collections of samples where visual inspection might be impractical, is therefore an important, and often far from trivial task, which remains the subject of active research [13].
A prominent example is Hartigan and Hartigan’s Dip Test [14], a widely used nonparametric method for assessing unimodality. Its statistic, known as the Dip, measures the maximal deviation between the empirical cumulative distribution function and the closest unimodal cumulative distribution, obtained via a linear envelope construction. By relying solely on order statistics and shape constraints rather than parametric assumptions, the Dip Test provides a simple and robust way to detect deviations from unimodality. Several alternative approaches can be considered, including Silverman’s kernel density-based test [15], excess mass tests [16, 17], and mode-tree methods [18]. While these methods offer complementary perspectives, they typically require tuning parameters, are computationally more demanding, or lack the conceptual immediacy of rank-based approaches. As a result, Dip Test-based methods have been adopted and continue to be widely used across a range of domains, from empirical studies of political behavior [19, 20, 21], to feature engineering [22] and deep learning-based clustering [23, 24, 25, 26, 27].
These methods, however, present a set of limitations that reduce their interpretability, especially in large-scale applications. First, critical values of the Dip statistic are strongly depend on sample size, thus requiring lookup tables or Monte Carlo simulation for assessment of significance. Second, the Dip statistic is not standardized, preventing meaningful comparisons across datasets of different sizes. Third, for large samples, the test rejects unimodality even for distributions that are in fact unimodal, since the null distribution of the Dip statistic narrows as the sample size increases. These issues complicate the use of Dip Test-based methods when distributional comparisons are performed across many subpopulations or time points, or when the goal is not only to test a deviation from unimodality, but to quantify its extent.
In this paper we introduce Z-Dip, a standardized measure of multimodality designed to address these challenges. Rather than treating modality detection purely as a hypothesis-testing problem, we rework the Dip statistic as a standardized score measuring the strength of deviation from unimodality. Specifically, we treat the Dip statistic as a random variable under the null hypothesis of unimodality and standardize its observed value using the mean and variance of its null distribution. This transformation converts the Dip from a sample-size dependent test statistic into a scale-independent measure of multimodality that is directly comparable across datasets and sample sizes. We show that the resulting score has a stable null distribution across a wide range of sample sizes, enabling the use of a single universal threshold for multimodality detection while preserving the original test’s computational complexity and parameter-free formulation. In this way, Z-Dip extends the Dip test from a binary hypothesis test to a continuous, standardized measure of multimodality, while retaining the original decision rule as a simple threshold on the standardized score.
We validate the Z-Dip on both simulated and empirical data. Simulated tests demonstrate that the Z-Dip reproduces Dip Test decisions across a number of sample sizes while providing a more interpretable scale. An empirical analysis of more than 88,000 opinion distributions shows practically perfect agreement between the Z-Dip threshold and the classical Dip Test. Finally, we present an effective downsampling-based correction that mitigates residual sample-size effects in very large datasets. An open-source R implementation of the Z-Dip is made available at https://github.com/EdoardoDima/zdip.
The rest of the paper is organized as follows: Section 2 introduces the theoretical background of Hartigans’ Dip Test and outlines its limitations. We then formally introduce the Z-Dip in Section 3, and show that the standardized formulation admits a universal significance threshold that closely reproduces the decisions of the classical Dip Test without requiring sample-size-specific calibration. We provide results of its calibration and validation in Section 4, using both simulated and empirical data, and, finally, we discuss our proposed correction for the distortion of these measures at large sample sizes in Section 5. Considerations regarding our findings and results can be found in Section 6.
@@ 2 Background
2.1 Hartigans’ Dip Test
The Dip Test, introduced by Hartigan and Hartigan in 1985 [14], is a classical nonparametric procedure for testing whether a distribution is unimodal. Given a sample $X_{1},\ldots,X_{N}$ drawn from an unknown distribution with Cumulative Distribution Function (CDF) $F$, the test evaluates the null hypothesis that $F$ is unimodal against the alternative that it is not.
A distribution is unimodal if its density has a single local maximum. Equivalently, its CDF is convex up to the mode and concave thereafter. Let $\mathcal{U}$ denote the class of unimodal distribution functions on $\mathbb{R}$. The Dip Test measures how far the empirical cumulative distribution function $F_{N}(x)$ deviates from this class.
Formally, the Dip statistic is defined as
$\mathrm{Dip}(F_{N})=\inf_{G\in\mathcal{U}}\sup_{x}|F_{N}(x)-G(x)|,$
that is, the smallest Kolmogorov–Smirnov distance between the empirical distribution function and the class of unimodal CDFs.
Hartigan and Hartigan showed that this quantity can be computed efficiently by exploiting the shape constraints implied by unimodality. In particular, the algorithm constructs two envelopes around the Empirical Cumulative Distribution Function (ECDF): the greatest convex minorant (GCM), defined as the largest convex function lying below $F_{N}$, and the least concave majorant (LCM), defined as the smallest concave function lying above $F_{N}$. These envelopes define candidate unimodal CDFs by joining segments where the slopes satisfy the compatibility condition
$\nabla\mathrm{GCM}\leq\nabla\mathrm{LCM}.$
The Dip statistic corresponds to the maximum vertical distance between the ECDF and the closest admissible unimodal CDF obtained through this construction. In more intuitive terms, the Dip statistic quantifies the smallest perturbation required to transform the empirical distribution into one that satisfies the unimodality constraint, where larger values therefore indicate stronger departures from unimodality. To assess statistical significance, the observed Dip value is compared to its sampling distribution under the null hypothesis of unimodality. Hartigan and Hartigan proposed using the uniform distribution on $[0,1]$ as a reference null, as it represents a least-favorable unimodal distribution. In practice, significance can be evaluated either through Monte Carlo simulation, by generating samples of size $N$ from a uniform distribution and computing their Dip statistics, or by consulting tabulated critical values provided in [14]. If the observed Dip exceeds the corresponding critical value, unimodality is rejected.
A key strength of the Dip Test lies in its simplicity: it is fully nonparametric and requires no user-specified smoothing parameters or model assumptions. Unlike kernel-based or mixture-based methods, it can be applied directly to empirical data without tuning or model selection, which has contributed to its wide adoption across many fields.
@@ 2.2 Limitations of the classical Dip Test
    Dependence on sample size. The null distribution of the Dip statistic varies strongly with sample size $N$. As $N$ increases, the Dip values expected under unimodality decrease systematically, since larger samples provide finer-grained approximations to the true underlying distribution, which has a flat density, and thus a linear CDF. Consequently, each sample size requires its own set of critical values or a separate Monte Carlo calibration. This dependence complicates both interpretation and automation, as results from samples of different sizes cannot be directly compared.
@@
    Lack of standardization. The Dip statistic itself is not standardized. Its raw magnitude depends on the scale of sampling variability and on $N$, making it unsuitable for quantitative comparison across datasets. For instance, a Dip of $0.1$ may signal multimodality in a small sample, but be negligible in a large one. Without a standardized scale, the test can only yield binary decisions (reject or not reject unimodality), rather than interpretable effect sizes that quantify “how multimodal” a distribution is. However, the Dip statistic already encapsulates this information: it measures the strength of deviation from unimodality, but this signal is obscured by sample-size dependence.
@@
    Inconsistencies at large sample sizes. As already mentioned, for large $N$, the Dip Test becomes increasingly sensitive to even minor deviations from the ideal unimodal shape. Because the null distribution of the Dip shrinks toward zero as $N$ grows, the smallest irregularities, such as sampling noise or slight skewness, are sufficient to reject unimodality. In practice, this produces false positives in large samples: distributions that are essentially unimodal, but statistically classified as multimodal. This issue means that the Dip Test’s binary decision rule becomes less informative with increasing $N$, effectively turning every large sample into a “significant” case. As a result, the test loses its ability to distinguish meaningful multimodality (e.g., distinct peaks) from trivial numerical deviations.
Together, these limitations point to the need for a standardized, scale-independent reformulation of the Dip Test that preserves its parameter-free simplicity while allowing meaningful comparison across sample sizes, maintaining interpretability at large $N$, and eliminating the need for case-specific calibration. The next section introduces the Z-Dip, which addresses these issues by standardizing the Dip relative to its null distribution.
@@ 3.1 Definition of the Z-Dip Statistic
We have already introduced how the behavior of the classical Dip statistic is strongly influenced by sample size: as $N$ increases, the empirical cumulative distribution function converges uniformly to the underlying distribution, and the null distribution of the Dip contracts toward zero. Essentially, the Dip under unimodality scales with the resolution of the ECDF, which increases with $N$, and, in consequence, the mean and variance of the null Dip decrease as $N$ grows.
@@
To address this, we propose a standardized version of Hartigan’s Dip statistic, which we call the Z-Dip. For an observed dataset of size $N$, let $\text{Dip}(F_{N})$ denote its Dip statistic’s value. We define
$Z\text{-}Dip=Z_{N}=\frac{\text{Dip}(F_{N})-\mu_{N}}{\sigma_{N}},$ (1)
@@
where $\mu_{N}$ and $\sigma_{N}$ denote the mean and standard deviation of Dip values under the null distribution of $N$-sized samples drawn from the uniform distribution on $[0,1]$, consistently with the original formulation of the Dip test. Standardization, therefore, removes the dominant sample-size effect. If $\mu_{N}$ and $\sigma_{N}$ equal the true null moments, then under unimodality the standardized statistic satisfies
$\mathbb{E}_{0}[Z_{N}]=0,\qquad\mathrm{Var}_{0}(Z_{N})=1,$
so that the null distribution is centered and variance-stabilized across $N$.
Although a full asymptotic characterization of the Dip statistic is beyond the scope of this work, empirical evidence suggests that the standardized statistic exhibits approximate distributional stability across a wide range of sample sizes under unimodality, with minor finite-sample deviations from normality, also explaining why a universal decision threshold can be derived and employed across a wide range of sample sizes. In summary, the score proposed in Formula (1) quantifies how strongly the observed distribution deviates from unimodality, in terms of standard deviation relative to the null expectation. Because it rescales the Dip relative to its null, the Z-Dip provides a scale-independent, comparable measure across different sample sizes. Furthermore, while the Z-Dip is inherently unbounded, in contrast to the classical Dip statistic, monotone transformations may be applied if a bounded scale is preferred (Fig. S1 of the Supplementary Information (SI)).
3.2 Asymptotic behavior
The behavior of the Z-Dip can be understood by examining how the magnitude of sampling fluctuations in the empirical distribution scales with the sample size.
Let $X_{1},\dots,X_{N}$ be independent samples drawn from a distribution with cumulative distribution function $F$. The empirical distribution
$F_{N}(x)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{X_{i}\leq x\},\qquad x\in\mathbb{R}$
converges uniformly to the true distribution. In particular, the Dvoretzky–Kiefer–Wolfowitz inequality states that for any $\varepsilon>0$
$\Pr\!\left(\sup_{x\in\mathbb{R}}|F_{N}(x)-F(x)|>\varepsilon\right)\leq 2e^{-2N\varepsilon^{2}}.$ (2)
It can also be shown that, by fixing a target error probability $\alpha$ such that $\alpha=\Pr\!\left(\sup_{x\in\mathbb{R}}|F_{N}(x)-F(x)|>\varepsilon\right)=2e^{-2N\varepsilon^{2}}$, solving for the deviation between the ECDF and the CDF returns
$\varepsilon=\sqrt{\frac{\log(\frac{2}{\alpha})}{2N}}=\sqrt{\frac{\log(\frac{2}{\alpha})}{2}}\cdot N^{-1/2}$
implying that the maximum vertical distance between the empirical and true CDF shrinks at rate $O_{p}(N^{-1/2})$.
Scaling under unimodality
Given these considerations, first consider the case where the underlying distribution $F$ is unimodal. Since the Dip statistic measures the minimal shift required to transform the empirical distribution into one that satisfies the unimodality constraint:
$\text{Dip}(F_{N})=\inf_{G\in U}\sup_{x\in\mathbb{R}}|F_{N}(x)-G(x)|,$
where $U$ denotes the set of unimodal cumulative distribution functions identified by the algorithm, it follows that, when the underlying distribution $F$ is unimodal, the true distribution itself belongs to the admissible class $U$. Consequently,
$\text{Dip}(F_{N})\leq\sup_{x\in\mathbb{R}}|F_{N}(x)-F(x)|.$
Since the empirical deviation from the CDF scales as $O_{p}(N^{-1/2})$, the Dip statistic must also shrink at least at this rate. This was already shown by Hartigan and Hartigan in their original paper, where it was noted how $\mathrm{Dip}(F_{N})\rightarrow 0$. At the same time, given the Z-Dip formulation:
$Z_{N}=\frac{\text{Dip}(F_{N})-\mu_{N}}{\sigma_{N}},$
the standard deviation of the Dip statistic under the null distribution is driven by the same process, resulting in both the numerator and denominator to scale at the same rate in $N$, leading to a Z-Dip distribution that remains stable around 0 across sample sizes.
Scaling under multimodality
If instead the underlying distribution $F$ is not unimodal, it will not be part of the set $U$ of candidate unimodal distributions.
Let us consider $\tilde{G}\in U$ and $\bar{x}\in\mathbb{R}$ such that
$\text{Dip}(F_{N})=|F_{N}(\bar{x})-\tilde{G}(\bar{x})|.$
There exists a positive constant $c=c(\tilde{G},\bar{x})$ such that
$c=|F(\bar{x})-\tilde{G}(\bar{x})|,$
and one simply has:
${\rm{Dip}}(F_{N})=|F_{N}(\bar{x})-\tilde{G}(\bar{x})|=|F_{N}(\bar{x})-F(\bar{x})+F(\bar{x})-\tilde{G}(\bar{x})|.$
Since the empirical CDF converges to $F$, the right-hand side of the inequality asymptotically vanishes, implying:
$\mathrm{Dip}(F_{N})\rightarrow c,$
meaning that, for distributions that deviate from unimodality, the Dip statistic converges to a positive constant $c$ reflecting the distance between the true distribution and the class of unimodal distributions. Substituting this property into our statistic and being $\mu_{N}\to 0$ as $N\to\infty$, we have
$Z_{N}=\frac{\mathrm{Dip}(F_{N})-\mu_{N}}{\sigma_{N}}\approx\frac{c}{\sigma_{N}}\quad\text{as }N\to\infty.$
We observe that the numerator converges asymptotically to a positive constant, while the standard deviation $\sigma_{N}$ continues to shrink proportionally to $N^{-1/2}$ by the Chebychev inequality and according with Formula (2). Consequently,
$Z_{N}\sim\frac{c}{N^{-1/2}}\propto N^{1/2},$
explaining its empirical power-law behavior for multimodal cases.
Empirical validation of asymptotic assumptions
These asymptotic considerations are also supported by empirical simulation results, as shown in Figure 1. Under the proposed asymptotic regime described above, if the underlying distribution deviates from unimodality, then the Dip statistic converges to a positive constant $c$, while the null standard deviation scales as $N^{-1/2}$; thus,
$\frac{Z_{N}}{N^{1/2}}\sim c,$
This behavior is clearly observed in Figure 1, where, for unimodal distributions (uniform and gaussian), the rescaled statistic $Z_{N}/N^{1/2}$ converges to zero, indicating that no persistent structural deviation is present and the observed values are driven entirely by sampling noise. In contrast, for multimodal distributions $Z_{N}/N^{1/2}$ rapidly stabilizes to a positive constant, where the different plateau levels correspond to different values of $c$, capturing the strength of the underlying multimodal structure. Moreover, the fact that $Z_{N}/N^{1/2}$ converges to a constant confirms that $Z_{N}$ grows proportionally to $N^{1/2}$.
Figure 1: Scaling of Z-Dip across modality regimes. Under unimodality, the rescaled statistic $Z_{N}/N^{1/2}$ converges to zero. Under multimodality, it stabilizes to a positive constant, consistent with the theoretical scaling $Z_{N}\propto N^{1/2}$. Plateau differences indicate varying strengths of multimodal structure.